The difference of two numbers is 4. Two times the smaller is 10 more than the larger. Find the two numbers.
Answer by Cromlix(4381)

Hi,
Make the larger number = x
Make the smaller number = y
x - y = 4..........Eq(1)
x + 10 = 2y........Eq(2)
Rearrange Eq(1)
x - y = 4
x = y + 4
Substitute x = y + 4 into Eq(2)
x + 10 = 2y
y + 4 + 10 = 2y
Collect like terms
2y - y = 14
y = 14
Substitute y = 14 into Eq(1)
x - y = 4
x - 14 = 4
x = 18
Larger number = 18
Smaller number = 14
Hope this helps :-)